MEASUREMENT OF THE ELECTRON BINDING ENERGY OF SIZE SELECTED IBr^{−}(CO_{2})_{n} CLUSTERS THROUGH ANION PHOTOELECTRON IMAGING

\underline{RYAN M. D. CALVI}, DJANGO H. ANDREWS, and W. CARL LINEBERGER, JILA and Department of Chemistry and Biochemistry, University of Colorado and National Institute of Standards and Technology, Campus Box 440, University of Colorado, Boulder, Colorado 80309-0440.

We report the measurement of the electron binding energies of size selected IBr^{−}(CO_{2})_{n}, n<10 cluster anions. The data are interpreted using ab initio calculations and Franck-Condon simulations to assist in the assignment of the photoelectron spectra. The IBr(CO_{2})_{n} anion are formed by a 1 kV electron impact on dilute IBr and CO_{2} in a pulsed supersonic expansion. Time-of-flight mass spectrometry allows a size-selected cluster anion to interact with a pulsed ultraviolet laser radiation tunable between 210 and 320 nm. The photodetached electrons are energy analyzed using a photoelectron imaging spectrometer operating in the velocity map imaging mode. The spectrometer determines electron kinetic energy, with an energy resolution of 2.5 percent. Trends in electron binding energy as a function of size will be discussed in regard to the trend that develops upon the addition of each cluster, and compared to previous photoelectron spectroscopy of I_{2}^{−}(CO_{2})_{n} clusters. Supported by NSF and AFOSR.